Explain why the following reactions are not allowed:
(a) 60 28 Ni (α, p) 62 29 Cu
(b) 27 13 Al (n, n) 28 13 Al
(c) 39 19 K (p, α) 36 17 Cl
Answer #1

har + He — 2 Qu + H Total maks no.=co+4 Total make no. = 62+1=63 =64 Total atomie no=28=211 Total atomie ho. = 29+1 = 30